# Andreas Petrou-Zeniou

### v3

@@ Appendix C Proofs of Other Theoretical Results
$\{AU\leq b\}$ =$=\{A(c(B^{\prime}U)+Z)\leq b\}$
=$=\{(Ac)(B^{\prime}U)\leq b-AZ\}$
yielding a set of linear constraints on $B^{\prime}U$, when conditioning on $Z$. Thus, because $Z$ is independent of $B^{\prime}U$, we find that $B^{\prime}U$, conditional on the selection event $AU\leq b$ and sufficient statistic $Z$, [^8] Here, $Z$ can be thought of as a sufficient statistic for the nuisance parameters in our model. is distributed according to a multivariate normal with mean $\mu_{B}:=B^{\prime}\mu$ and variance-covariance $\Sigma_{B}:=B^{\prime}\Sigma B$, truncated to the polyhedron $\mathcal{O}:=\{x;(Ac)x\leq b-AZ\}$.
@@
Proof of Proposition 4.1. First, we notice that $\left\{\mu_{Y,\jmath}\right\}_{\jmath\in\hat{J}_{R}}\in CS_{1-\alpha}^{c}$ if and only if our test $\phi\left(\ \cdot\ ;(i_{l})_{l\in R},z,(\mu_{Y,i_{l}})_{l\in R}\right)$ fails to reject. Because this test is a valid test at level $\alpha$, (5) holds.
@@
$P_{\mu,\Sigma}\left(\bigcup_{\jmath\in\hat{J}_{R}}\left\{\frac{|\xi_{Y,\jmath}|}{\sqrt{\Sigma_{Y,\jmath\jmath}}}>\rho_{1-\alpha+\beta}(L,U)\right\}\right)$
$\leq P_{\mu,\Sigma}\left(\bigcup_{\jmath\in\hat{J}_{R}}\left\{\frac{|\xi_{Y,\jmath}|}{\sqrt{\Sigma_{Y,\jmath\jmath}}}>\rho_{1-\alpha+\beta}(L,U)\right\}\cap B\right)+P(B^{c})$
$\leq P_{\mu,\Sigma}\left(\bigcup_{\jmath\in\hat{J}_{R}}\left\{\frac{|\xi_{Y,\jmath}|}{\sqrt{\Sigma_{Y,\jmath\jmath}}}>\rho_{1-\alpha+\beta}(\Delta,\Delta)\right\}\cap B\right)+\beta$
$\leq P_{\mu,\Sigma}\left(\bigcup_{\jmath\in\hat{J}_{R}}\left\{\frac{|\xi_{Y,\jmath}|}{\sqrt{\Sigma_{Y,\jmath\jmath}}}>\rho_{1-\alpha+\beta}(\Delta,\Delta)\right\}\right)+\beta$
$\leq P_{\mu,\Sigma}\left(\max_{j\in J}\frac{|\xi_{Y,j}|}{\sqrt{\Sigma_{Y,jj}}}\mathbb{1}\left(\sum_{j^{\prime}\in J}\mathbb{1}\left(\xi_{X,j}\geq\xi_{X,j^{\prime}}+\Delta_{j^{\prime},j}\right)\in R\right)>\rho_{1-\alpha+\beta}(\Delta,\Delta)\right)+\beta$
@@
$P_{\mu,\Sigma}\left(\bigcap_{\jmath\in\hat{J}_{R^{\prime}}}\left\{\frac{|\xi_{Y,\jmath}|}{\sqrt{\Sigma_{Y,\jmath\jmath}}}\leq\rho_{1-\alpha+\beta}(L,U)\right\}\right)\geq 1-\alpha$
@@
We now provide a proof of proposition 5.2, which proceeds much as the proof of proposition 5.1.
@@
Proof of Proposition 5.2. We divide our proof into cases. In the first case, we have $\rho_{1-\alpha+\beta}(\Delta,\Delta)\leq\bar{c}_{1-\alpha}$. In the second case, we have $\rho_{1-\alpha+\beta}(\Delta,\Delta)>\bar{c}_{1-\alpha}$. The proof of validity in the first case proceeds exactly as in 5.1, so we omit details and focus on proving validity in the second case. As before, we take $B:=\{L\leq\Delta\leq U\}$. In the second case, we have the following:
$P_{\mu,\Sigma}\left(\mu_{Y,\jmath}\in CS_{1-\alpha}^{TS2}\quad\text{for all }\jmath\in\hat{J}_{R}\right)$
$\geq P_{\mu,\Sigma}\left(\left\{\mu_{Y,\jmath}\in CS_{1-\alpha}^{TS2}\quad\text{for all }\jmath\in\hat{J}_{R}\right\}\cap B\right)$
$\geq P_{\mu,\Sigma}\left(\left\{\mu_{Y,\jmath}\in\left[Y_{\jmath}\pm\left(\rho_{\alpha,\beta}(\Delta,\Delta)\wedge\bar{c}_{1-\alpha}\right)\sqrt{\Sigma_{Y,\jmath\jmath}}\right]\quad\text{for all }\jmath\in\hat{J}_{R}\right\}\cap B\right)$
$\geq P_{\mu,\Sigma}\left(\left\{\mu_{Y,\jmath}\in\left[Y_{\jmath}\pm\bar{c}_{1-\alpha}\sqrt{\Sigma_{Y,\jmath\jmath}}\right]\quad\text{for all }\jmath\in\hat{J}_{R}\right\}\cap B\right)$
@@
Finally, to prove proposition 5.3, we impose a new assumption. In particular, we impose the following pointwise rate of convergence on our variance-covariance estimator $\hat{\Sigma}^{n}$, which is necessary in proving almost sure convergence of variance-covariance in the proof for proposition 5.3.
Assumption C.1.
We assume that, for any $P$ in $\mathcal{P}$, there exists $\delta>0$ such that, for any $\varepsilon>0$:
$Pr_{P}\left(\left\|\hat{\Sigma}^{n}(P)-\Sigma(P)\right\|>\varepsilon\right)=o(1/n^{\delta})$
where our choice of matrix norm is again the max norm.
@@
Proof of Proposition 5.3. We begin by recalling that:
$L_{j^{\prime},j}^{n}$ $=\tilde{S}_{X,j^{\prime}}^{n}-\tilde{S}_{X,j}^{n}-d_{1-\beta/2}(\hat{\Sigma}^{n})\sqrt{\frac{\hat{\text{var}}^{n}_{j^{\prime}j}}{n}}$
$U_{j^{\prime},j}^{n}$ $=\tilde{S}_{X,j^{\prime}}^{n}-\tilde{S}_{X,j}^{n}+d_{1-\beta/2}(\hat{\Sigma}^{n})\sqrt{\frac{\hat{\text{var}}^{n}_{j^{\prime}j}}{n}}$
We note that, by assumption B.2, for fixed $P$ in $\mathcal{P}$, $\left\|\hat{\Sigma}^{n}-\Sigma(P)\right\|=O(1)$ almost surely. Consequently, by continuity, $d_{1-\beta/2}(\hat{\Sigma}^{n})=O(1)$ almost surely, while the vector $\hat{\text{var}}^{n}/n=o(1)$ almost surely. Thus:
$\begin{pmatrix}L_{j^{\prime},j}^{n}\\ U_{j^{\prime},j}^{n}\end{pmatrix}\xrightarrow[]{a.s.}\begin{pmatrix}\Delta_{j^{\prime},j}(P)\\ \Delta_{j^{\prime},j}(P)\end{pmatrix}$ (36)
@@
We define $\delta:=\min_{j\in J_{R}(P),i\not\in J_{R}(P)}\left|\mu_{X,j}-\mu_{X,i}\right|$. Naturally, because $J_{R}(P)$ is a proper subset of $J$, $\delta>0$. It follows from (36) that the event $B:=\left\{\left|L^{n}-\Delta\right|,\ \left|U^{n}-\Delta\right|\leq\delta/6\ e.v.\right\}$ holds almost surely. That is, $P(B)=1$. We seek to study the convergence of $\rho_{1-\alpha+\beta}\left(L^{n},U^{n};\hat{\Sigma}^{n}/n\right)$ on $B$. We now seek to show that:
$\mathbb{1}\left(\left[\sum_{j^{\prime}\in J}\mathbb{1}\left(\xi_{X,j}^{n}\geq\xi_{X,j^{\prime}}^{n}+U_{j^{\prime},j}^{n}\right),\sum_{j^{\prime}\in J}\mathbb{1}\left(\xi_{X,j}^{n}\geq\xi_{X,j^{\prime}}^{n}+L_{j^{\prime},j}^{n}\right)\right]\cap R\neq\emptyset\right)\to\mathbb{1}\left(j\in J_{R}(P)\right)$ (37)
in probability, conditional on $L^{n}$ and $U^{n}$. We note that $\xi_{W}^{n}$ are drawn from a Gaussian with mean zero and variance-covariance $\hat{\Sigma}^{n}/n$, and are defined to be independent from the $\tilde{X}_{j}$ and $\tilde{Y}$. Rigorously, we want that:
$Pr_{P}\left(\mathbb{1}\left(\left[\sum_{j^{\prime}\in J}\mathbb{1}\left(\xi_{X,j}^{n}\geq\xi_{X,j^{\prime}}^{n}+U_{j^{\prime},j}^{n}\right),\sum_{j^{\prime}\in J}\mathbb{1}\left(\xi_{X,j}^{n}\geq\xi_{X,j^{\prime}}^{n}+L_{j^{\prime},j}^{n}\right)\right]\cap R\neq\emptyset\right)\neq\mathbb{1}\left(j\in J_{R}(P)\right)|L^{n},U^{n}\right)$ (38)
is $o(1)$ as $n$ approaches infinity. Consequently, we can show that, [^9] This result can be shown by recalling the Kolmogorov’s definition of conditional probability, and subsequently applying Fubini’s theorem. on the event $\{L^{n}=l^{n},U^{n}=u^{n}\}$,
$Pr_{P}\left(\mathbb{1}\left(\left[\sum_{j^{\prime}\in J}\mathbb{1}\left(\xi_{X,j}^{n}\geq\xi_{X,j^{\prime}}^{n}+U_{j^{\prime},j}^{n}\right),\sum_{j^{\prime}\in J}\mathbb{1}\left(\xi_{X,j}^{n}\geq\xi_{X,j^{\prime}}^{n}+L_{j^{\prime},j}^{n}\right)\right]\cap R\neq\emptyset\right)\neq\mathbb{1}\left(j\in J_{R}(P)\right)|L^{n},U^{n}\right)$
$=Pr_{P}\left(\mathbb{1}\left(\left[\sum_{j^{\prime}\in J}\mathbb{1}\left(\xi_{X,j}^{n}\geq\xi_{X,j^{\prime}}^{n}+u_{j^{\prime},j}^{n}\right),\sum_{j^{\prime}\in J}\mathbb{1}\left(\xi_{X,j}^{n}\geq\xi_{X,j^{\prime}}^{n}+l_{j^{\prime},j}^{n}\right)\right]\cap R\neq\emptyset\right)\neq\mathbb{1}\left(j\in J_{R}(P)\right)\right)$ (39)
Now, we define $B_{n}$ to be correspond to the event that all $\xi_{X,j}^{n}$ are within $\delta/6$ of zero. We note that zero is simply the probability limit of the $\xi_{X,j}^{n}$, implying that $P(B_{n})$ approaches one as $n$ approaches infinity. It is clear that, taking $l^{n}$ and $u^{n}$ arbitrarily such that $|l^{n}-\Delta|\leq\delta/6$ and $|u^{n}-\Delta|\leq\delta/6$:
@@
$\lim_{n\to\infty}Pr_{P}\left(\mathbb{1}\left(\left[\sum_{j^{\prime}\in J}\mathbb{1}\left(\xi_{X,j}^{n}\geq\xi_{X,j^{\prime}}^{n}+u_{j^{\prime},j}^{n}\right),\sum_{j^{\prime}\in J}\mathbb{1}\left(\xi_{X,j}^{n}\geq\xi_{X,j^{\prime}}^{n}+l_{j^{\prime},j}^{n}\right)\right]\cap R\neq\emptyset\right)\neq\mathbb{1}\left(j\in J_{R}(P)\right)\right)$
$=\lim_{n\to\infty}Pr_{P}\left(\left\{\mathbb{1}\left(\left[\sum_{j^{\prime}\in J}\mathbb{1}\left(\xi_{X,j}^{n}\geq\xi_{X,j^{\prime}}^{n}+u_{j^{\prime},j}^{n}\right),\sum_{j^{\prime}\in J}\mathbb{1}\left(\xi_{X,j}^{n}\geq\xi_{X,j^{\prime}}^{n}+l_{j^{\prime},j}^{n}\right)\right]\cap R\neq\emptyset\right)\neq\mathbb{1}\left(j\in J_{R}(P)\right)\right\}\cap B_{n}\right)$
$=\lim_{n\to\infty}Pr_{P}(\emptyset)=0$
almost surely. Thus, (38) holds. Moreover, leveraging the above convergence in conditional probability result, we note that, by the continuous mapping theorem, the right hand side of (12) converges to:
$\max_{j\in J_{R}(P)}\frac{\left|\xi_{Y,j}\right|}{\sqrt{\Sigma_{Y,jj}}}$
in distribution, given $L^{n}=l^{n}$ and $U^{n}=u^{n}$ chosen above. We note that $\xi_{Y}$ is distributed according to a multivariate Gaussian with mean zero and variance-covariance $\Sigma_{Y}(P)$. Thus, $\rho_{\alpha,\beta}\left(L^{n},U^{n};\hat{\Sigma}^{n}\right)$ converges to $c_{1-\alpha+\beta}\left(J_{R}(P)\right)$ as $n$ grows large, on $B$. It naturally follows that, for sufficiently small $\beta$, $c_{1-\alpha+\beta}\left(J_{R}(P)\right)<c_{1-\alpha}\left(J\right)$, as $c_{\gamma}(J_{c})$ is continuous in $\gamma$ for any $J_{c}$ in $2^{J}$.
It now remains to prove that equation 20 holds. To do so, we pick some small, positive value of $h$ such that we can construct a function $\mathbb{1}_{h}$ satisfying:
$\mathbb{1}(x>y)\geq\mathbb{1}_{h}(x,y)\geq\mathbb{1}(x>y+h)$
and such that $\mathbb{1}_{h}$ is continuous in $x$ and $y$. For $k$ in $R$
We also let $M_{n}$ be the event where $\hat{J}_{R;n}=J_{R}(P)$. By previous arguments, we know that the set $\{M_{n}\ e.v.\}$ holds almost surely. We can now make the following calculations:
$\lim_{n\to\infty}\mathbb{E}_{P}\left(\mathbb{1}_{h}\left(\max_{\jmath\in\hat{J}_{R;n}}\frac{\left|\tilde{S}_{Y,\jmath}^{n}-\mu_{Y,\jmath}\left(P\right)\right|}{\sqrt{\hat{\Sigma}_{Y,\jmath\jmath}^{n}/n}},\rho_{\alpha,\beta}(L^{n},U^{n},\hat{\Sigma}^{n}/n)\right)\right)$
$=\mathbb{E}_{P}\left(\lim_{n\to\infty}\mathbb{1}_{h}\left(\max_{\jmath\in\hat{J}_{R;n}}\frac{\left|\tilde{S}_{Y,\jmath}^{n}-\mu_{Y,\jmath}\left(P\right)\right|}{\sqrt{\hat{\Sigma}_{Y,\jmath\jmath}^{n}/n}},\rho_{\alpha,\beta}(L^{n},U^{n},\hat{\Sigma}^{n}/n)\right)\right)$
$=\mathbb{E}_{P}\left(\lim_{n\to\infty}\mathbb{1}_{h}\left(\max_{j\in J_{R}(P)}\frac{\left|\tilde{S}_{Y,j}^{n}-\mu_{Y,j}\left(P\right)\right|}{\sqrt{\hat{\Sigma}_{Y,jj}^{n}/n}},c_{1-\alpha+\beta}\left(J_{R}(P)\right)\right)\right)$
$\geq\mathbb{E}_{P}\left(\lim_{n\to\infty}\mathbb{1}_{h}\left(\max_{j\in J_{R}(P)}\frac{\left|\tilde{S}_{Y,j}^{n}-\mu_{Y,j}\left(P\right)\right|}{\sqrt{\hat{\Sigma}_{Y,jj}^{n}/n}},c_{1-\alpha}\left(J\right)\right)\right)$
$=\lim_{n\to\infty}\mathbb{E}_{P}\left(\mathbb{1}_{h}\left(\max_{\jmath\in\hat{J}_{R;n}}\frac{\left|\tilde{S}_{Y,\jmath}^{n}-\mu_{Y,\jmath}\left(P\right)\right|}{\sqrt{\hat{\Sigma}_{Y,\jmath\jmath}^{n}/n}},c_{1-\alpha}\left(J\right)\right)\right)$
That
$\lim_{n\to\infty}\mathbb{E}_{P}\left(\mathbb{1}_{h}\left(\max_{\jmath\in\hat{J}_{R;n}}\frac{\left|\tilde{S}_{Y,\jmath}^{n}-\mu_{Y,\jmath}\left(P\right)\right|}{\sqrt{\hat{\Sigma}_{Y,\jmath\jmath}^{n}/n}},\rho_{\alpha,\beta}(L^{n},U^{n},\hat{\Sigma}^{n}/n)\right)\right)$
exists holds by the central limit theorem, [^10] The central limit theorem applies pointwise to $P$, since it applies uniformly over $\mathcal{P}$, under assumption B.1. continuous mapping theorem, and portmanteau lemma. The same reasoning applies to show the existence of:
$\lim_{n\to\infty}\mathbb{E}_{P}\left(\mathbb{1}_{h}\left(\max_{\jmath\in\hat{J}_{R;n}}\frac{\left|\tilde{S}_{Y,\jmath}^{n}-\mu_{Y,\jmath}\left(P\right)\right|}{\sqrt{\hat{\Sigma}_{Y,\jmath\jmath}^{n}/n}},c_{1-\alpha}\left(J\right)\right)\right)$
Thus, with these existence checks in mind, we have that, for arbitrary $h$, it follows that:

### v4

@@ Appendix C Proofs of Other Theoretical Results
$\{AU\leq b\}$ $=\{A(c(B^{\prime}U)+Z)\leq b\}$
$=\{(Ac)(B^{\prime}U)\leq b-AZ\}$
yielding a set of linear constraints on $B^{\prime}U$, when conditioning on $Z$. Thus, because $Z$ is independent of $B^{\prime}U$, we find that $B^{\prime}U$, conditional on the selection event $AU\leq b$ and sufficient statistic ZZ, [^9] Here, $Z$ can be thought of as a sufficient statistic for the nuisance parameters in our model. is distributed according to a multivariate normal with mean $\mu_{B}:=B^{\prime}\mu$ and variance-covariance $\Sigma_{B}:=B^{\prime}\Sigma B$, truncated to the polyhedron $\mathcal{O}:=\{x;(Ac)x\leq b-AZ\}$.
@@
Proof of Proposition 4.1. First, we notice that $\left\{\mu_{Y,\hat{\jmath}}\right\}_{\hat{\jmath}\in\hat{J}_{R}}\in CS_{1-\alpha}^{c}$ if and only if our test $\phi\left(\ \cdot\ ;(i_{l})_{l\in R},z,(\mu_{Y,i_{l}})_{l\in R}\right)$ fails to reject. Because this test is a valid test at level $\alpha$, (16) holds.
@@
$P_{\mu,\Sigma}\left(\bigcup_{\hat{\jmath}\in\hat{J}_{R}}\left\{\frac{|\xi_{Y,\hat{\jmath}}|}{\sqrt{\Sigma_{Y,\hat{\jmath}\hat{\jmath}}}}>\rho_{1-\alpha+\beta}(L,U)\right\}\right)$
$\leq P_{\mu,\Sigma}\left(\bigcup_{\hat{\jmath}\in\hat{J}_{R}}\left\{\frac{|\xi_{Y,\hat{\jmath}}|}{\sqrt{\Sigma_{Y,\hat{\jmath}\hat{\jmath}}}}>\rho_{1-\alpha+\beta}(L,U)\right\}\cap B\right)+P(B^{c})$
$\leq P_{\mu,\Sigma}\left(\bigcup_{\hat{\jmath}\in\hat{J}_{R}}\left\{\frac{|\xi_{Y,\hat{\jmath}}|}{\sqrt{\Sigma_{Y,\hat{\jmath}\hat{\jmath}}}}>\rho_{1-\alpha+\beta}(\Delta,\Delta)\right\}\cap B\right)+\beta$
$\leq P_{\mu,\Sigma}\left(\bigcup_{\hat{\jmath}\in\hat{J}_{R}}\left\{\frac{|\xi_{Y,\hat{\jmath}}|}{\sqrt{\Sigma_{Y,\hat{\jmath}\hat{\jmath}}}}>\rho_{1-\alpha+\beta}(\Delta,\Delta)\right\}\right)+\beta$
$\leq P_{\mu,\Sigma}\left(\max_{j\in J}\frac{|\xi_{Y,j}|}{\sqrt{\Sigma_{Y,jj}}}\mathbb{1}\left(\sum_{j^{\prime}\in J}\mathbb{1}\left(\xi_{X,j}\geq\xi_{X,j^{\prime}}+\Delta_{j^{\prime}j}\right)\in R\right)>\rho_{1-\alpha+\beta}(\Delta,\Delta)\right)+\beta$
@@
$P_{\mu,\Sigma}\left(\bigcap_{\hat{\jmath}\in\hat{J}_{R^{\prime}}}\left\{\frac{|\xi_{Y,\hat{\jmath}}|}{\sqrt{\Sigma_{Y,\hat{\jmath}\hat{\jmath}}}}\leq\rho_{1-\alpha+\beta}(L,U)\right\}\right)\geq 1-\alpha$
@@
We now provide a proof of proposition A.1, which proceeds much as the proof of proposition 3.1.
@@
Proof of Proposition A.1. We divide our proof into cases. In the first case, we have $\rho_{1-\alpha+\beta}(\Delta,\Delta)\leq\bar{c}_{1-\alpha}$. In the second case, we have $\rho_{1-\alpha+\beta}(\Delta,\Delta)>\bar{c}_{1-\alpha}$. The proof of validity in the first case proceeds exactly as in proposition 3.1, so we omit details and focus on proving validity in the second case. As before, we take $B:=\{L\leq\Delta\leq U\}$. In the second case, we have the following:
$P_{\mu,\Sigma}\left(\mu_{Y,\hat{\jmath}}\in CS_{1-\alpha;\beta}^{TS2}\quad\text{for all }\hat{\jmath}\in\hat{J}_{R}\right)$
$\geq P_{\mu,\Sigma}\left(\left\{\mu_{Y,\hat{\jmath}}\in CS_{1-\alpha;\beta}^{TS2}\quad\text{for all }\hat{\jmath}\in\hat{J}_{R}\right\}\cap B\right)$
$\geq P_{\mu,\Sigma}\left(\left\{\mu_{Y,\hat{\jmath}}\in\left[Y_{\hat{\jmath}}\pm\left(\rho_{1-\alpha+\beta}(\Delta,\Delta)\wedge\bar{c}_{1-\alpha}\right)\sqrt{\Sigma_{Y,\hat{\jmath}\hat{\jmath}}}\right]\quad\text{for all }\hat{\jmath}\in\hat{J}_{R}\right\}\cap B\right)$
$\geq P_{\mu,\Sigma}\left(\left\{\mu_{Y,\hat{\jmath}}\in\left[Y_{\hat{\jmath}}\pm\bar{c}_{1-\alpha}\sqrt{\Sigma_{Y,\hat{\jmath}\hat{\jmath}}}\right]\quad\text{for all }\hat{\jmath}\in\hat{J}_{R}\right\}\cap B\right)$
@@
Proof of Proposition 4.2. Let us fix some arbitrary $\beta>0$. We begin by recalling that:
$L_{j^{\prime}j}^{n}$ $=\tilde{S}_{X,j^{\prime}}^{n}-\tilde{S}_{X,j}^{n}-d_{1-\beta/2}(\hat{\Sigma}^{n})\sqrt{\hat{\text{var}}^{n}_{j^{\prime}j}}$
$U_{j^{\prime}j}^{n}$ $=\tilde{S}_{X,j^{\prime}}^{n}-\tilde{S}_{X,j}^{n}+d_{1-\beta/2}(\hat{\Sigma}^{n})\sqrt{\hat{\text{var}}^{n}_{j^{\prime}j}}$
We note that, by the law of large numbers and by the continuous mapping theorem, for fixed $P$ in $\mathcal{P}$, $\left\|n\hat{\Sigma}^{n}-\Sigma(P)\right\|=o_{P}(1)$. We also know that, for any fixed $\Sigma$, $d_{1-\beta/2}(\Sigma)$ is $O(1)$, meaning that $d_{1-\beta/2}(\hat{\Sigma}^{n})$ is $O_{P}(1)$. Finally, recalling that $\hat{\text{var}}^{n}_{j^{\prime}j}=\hat{\Sigma}^{n}_{j^{\prime}j^{\prime}}+\hat{\Sigma}^{n}_{jj}-2\hat{\Sigma}^{n}_{j^{\prime}j}$, we notice that $\hat{\text{var}}^{n}_{j^{\prime}j}=o_{P}(1)$. Thus, by the weak law of large numbers and that fact that the above facts imply that $d_{1-\beta/2}(\hat{\Sigma}^{n})\sqrt{\hat{\text{var}}^{n}_{j^{\prime}j}}=o_{P}(1)$:
$\begin{pmatrix}L_{j^{\prime}j}^{n}\\ U_{j^{\prime}j}^{n}\end{pmatrix}\xrightarrow[]{P}\begin{pmatrix}\Delta_{j^{\prime}j}(P)\\ \Delta_{j^{\prime}j}(P)\end{pmatrix}$ (43)
@@
We define $\delta:=\min_{j\in J_{R}(P),i\not\in J_{R}(P)}\left|\mu_{X,j}-\mu_{X,i}\right|$. Naturally, because $J_{R}(P)$ is a proper subset of $J$, $\delta>0$. It follows from (43) that the event $A_{n}:=\left\{\left|L^{n}-\Delta\right|,\ \left|U^{n}-\Delta\right|\leq\delta/6\right\}$ satisfies $P(A_{n})\to 1$. We seek to study the behavior of $\rho_{1-\alpha+\beta}\left(L^{n},U^{n};\hat{\Sigma}^{n}\right)$ on $A_{n}$. We now seek to show that:
$\left(\mathbb{1}\left(\left[\sum_{j^{\prime}\in J}\mathbb{1}\left(\xi_{X,j}^{n}\geq\xi_{X,j^{\prime}}^{n}+U_{j^{\prime}j}^{n}\right),\sum_{j^{\prime}\in J}\mathbb{1}\left(\xi_{X,j}^{n}\geq\xi_{X,j^{\prime}}^{n}+L_{j^{\prime}j}^{n}\right)\right]\cap R\neq\emptyset\right)\right)_{j\in J}\leq\left(\mathbb{1}\left(j\in J_{R}(P)\right)\right)_{j\in J}$ (44)
with probability approaching one. In what follows, we will suppress the vector notation above. We let $l^{n}$ and $u^{n}$ be such that $\left|l^{n}-\Delta\right|,\ \left|u^{n}-\Delta\right|\leq\delta/6$. We are interested in the following probability:
$Pr_{P}\left(\mathbb{1}\left(\left[\sum_{j^{\prime}\in J}\mathbb{1}\left(\xi_{X,j}^{n}\geq\xi_{X,j^{\prime}}^{n}+u_{j^{\prime}j}^{n}\right),\sum_{j^{\prime}\in J}\mathbb{1}\left(\xi_{X,j}^{n}\geq\xi_{X,j^{\prime}}^{n}+l_{j^{\prime}j}^{n}\right)\right]\cap R\neq\emptyset\right)\leq\mathbb{1}\left(j\in J_{R}(P)\right)\right)$
where $\xi_{W}^{n}$ are taken to have a Gaussian law with mean zero and variance-covariance $\hat{\Sigma}^{n}/n$. Now, we define $B_{n}$ to be correspond to the event that all $\xi_{X,j}^{n}$ are within $\delta/6$ of zero. Formally, $B_{n}:=\{\max_{j\in J}|\xi_{X,j}^{n}|<\delta/6\}$. Since $\xi_{X,j}^{n}\xrightarrow{P}0$, it follows that $P(B_{n})$ approaches one as $n$ approaches infinity. For such $l^{n}$ and $u^{n}$ as above, it is clear that:
$\lim_{n\to\infty}Pr_{P}\left(\mathbb{1}\left(\left[\sum_{j^{\prime}\in J}\mathbb{1}\left(\xi_{X,j}^{n}\geq\xi_{X,j^{\prime}}^{n}+u_{j^{\prime}j}^{n}\right),\sum_{j^{\prime}\in J}\mathbb{1}\left(\xi_{X,j}^{n}\geq\xi_{X,j^{\prime}}^{n}+l_{j^{\prime}j}^{n}\right)\right]\cap R\neq\emptyset\right)\leq\mathbb{1}\left(j\in J_{R}(P)\right)\right)$
$=\lim_{n\to\infty}Pr_{P}\left(\left\{\mathbb{1}\left(\left[\sum_{j^{\prime}\in J}\mathbb{1}\left(\xi_{X,j}^{n}\geq\xi_{X,j^{\prime}}^{n}+u_{j^{\prime}j}^{n}\right),\sum_{j^{\prime}\in J}\mathbb{1}\left(\xi_{X,j}^{n}\geq\xi_{X,j^{\prime}}^{n}+l_{j^{\prime}j}^{n}\right)\right]\cap R\neq\emptyset\right)\leq\mathbb{1}\left(j\in J_{R}(P)\right)\right\}\cap B_{n}\right)$
$=1-\lim_{n\to\infty}Pr_{P}\left(\left\{\exists j\not\in J_{R}(P);\left[\sum_{j^{\prime}\in J}\mathbb{1}\left(\xi_{X,j}^{n}\geq\xi_{X,j^{\prime}}^{n}+u_{j^{\prime}j}^{n}\right),\sum_{j^{\prime}\in J}\mathbb{1}\left(\xi_{X,j}^{n}\geq\xi_{X,j^{\prime}}^{n}+l_{j^{\prime}j}^{n}\right)\right]\cap R\neq\emptyset\right\}\cap B_{n}\right)$
@@
$=1-\lim_{n\to\infty}Pr_{P}\left(\left\{\exists j\not\in J_{R}(P);\left[\sum_{j^{\prime}\in J}\mathbb{1}\left(0\geq\Delta_{j^{\prime}j}+\delta/2\right),\sum_{j^{\prime}\in J}\mathbb{1}\left(0\geq\Delta_{j^{\prime}j}-\delta/2\right)\right]\cap R\neq\emptyset\right\}\right)$
$=1-Pr_{P}(\emptyset)=1$
Where we notice that $j\in J_{R}(P)$ if and only if:
$\sum_{j^{\prime}\in J}\mathbb{1}\left(0\geq\Delta_{j^{\prime}j}\right)\in R$
Suppose that there exists $j^{\prime}\not\in J_{R}(P)$ such that there exists $r\in R$ satisfying:
$r\in\left[\sum_{j^{\prime\prime}\in J}\mathbb{1}\left(0\geq\Delta_{j^{\prime\prime}j^{\prime}}+\delta/2\right),\sum_{j^{\prime\prime}\in J}\mathbb{1}\left(0\geq\Delta_{j^{\prime\prime}j^{\prime}}-\delta/2\right)\right]$
Let $j_{r}(\mu_{X})$ be some element in $J_{R}(P)$ such that $r=\sum_{j^{\prime\prime}\in J}\mathbb{1}\left(0\geq\Delta_{j^{\prime\prime}j_{r}(\mu_{X})}\right)$. Assume, without loss of generality, that $j^{\prime}$ is the index in $J_{R}(P)^{c}$ corresponding to the largest value of $\mu_{X,j}$ such that $\mu_{X,j}<\mu_{X,j_{r}(\mu_{X})}$. Consequently, we have that:
$\sum_{j^{\prime\prime}\in J}\mathbb{1}\left(0\geq\Delta_{j^{\prime\prime}j^{\prime}}+\delta/2\right)\leq\sum_{j^{\prime\prime}\in J}\mathbb{1}\left(0\geq\Delta_{j^{\prime\prime}j_{r}(\mu_{X})}\right)\leq\sum_{j^{\prime\prime}\in J}\mathbb{1}\left(0\geq\Delta_{j^{\prime\prime}j^{\prime}}-\delta/2\right)$
Clearly, this inequality cannot occur, since $|\Delta_{j^{\prime}j_{r}(\mu_{X})}|\geq\delta$. Thus, the event:
$\left\{\exists j\not\in J_{R}(P);\left[\sum_{j^{\prime}\in J}\mathbb{1}\left(0\geq\Delta_{j^{\prime}j}-\delta/2\right),\sum_{j^{\prime}\in J}\mathbb{1}\left(0\geq\Delta_{j^{\prime}j}+\delta/2\right)\right]\cap R\neq\emptyset\right\}$
is indeed the empty set. Consequently, whenever $l^{n}$ and $u^{n}$ are as above, $\rho_{1-\alpha+\beta}(l^{n},u^{n};\hat{\Sigma}^{n})\leq c_{1-\alpha+\beta}(J_{R}(P))$, with probability approaching one. Similarly, because $P(A_{n})\to 1$ and because $\|n\hat{\Sigma}^{n}-\Sigma(P)\|=o_{p}(1)$, we obtain $\rho_{1-\alpha+\beta}(L^{n},U^{n};\hat{\Sigma}^{n})\leq c_{1-\alpha+\beta}(J_{R}(P))$. If $\beta$ is sufficiently small such that $c_{1-\alpha+\beta}(J_{R}(P))\leq c_{1-\alpha}(J)$, we get $\rho_{1-\alpha+\beta}(L^{n},U^{n};\hat{\Sigma}^{n})\leq c_{1-\alpha}(J)$. It follows that for any $\beta>0$ sufficiently small as above: